On November 1, 2015, Reid Corporation acquired bonds with a face value of $700,000 for $673,618.61. The bonds carry a stated rate of interest of 10%, were purchased to yield 11%, pay interest semiannually on April 30 and October 31, were purchased to be held to maturity, and are due October 31, 2020. On November 1, 2016, in contemplation of a major acquisition, the bonds were sold for $700,000. Reid is on a fiscal year accounting period ending October 31 and uses the effective interest method.

Required:

Prepare journal entries to record the purchase of the bonds, the interest receipts on April 30, 2016, and October 31, 2016, and the sale of the bonds.
Answer #1
Date Account Debit Credit
Nov 1, 2015 Investment in bonds 6,73,618.61
Cash 6,73,618.61
[Entry to record purchase of the bonds]
April 30, 2016 Cash      35,000.00
Interest revenue      37,049.02
Investment in bonds        2,049.02
[Entry to record interest payment]
Oct 31, 2016 Cash      35,000.00
Interest revenue      37,161.72
Investment in bonds        2,161.72
[Entry to record interest payment]
Nov 1, 2016 Cash 7,00,000.00
Investment in bonds 6,77,829.35
Gain on sale of bonds      22,170.65
[Entry to record sale of bonds]

Interest and carrying value workings:

Date Beginning investment value Cash interest Interest revenue Increase in investment value Ending investment value
A B= 700,000 X 10%X 6/12 C= A X 11% X 6/12 D = C-B E= A+D
Oct-15                       6,73,618.61
Apr-16                             6,73,618.61                           35,000.00               37,049.02                                     2,049.02                       6,75,667.63
Oct-16                             6,75,667.63                           35,000.00               37,161.72                                     2,161.72                       6,77,829.35